A person is traveling to Chicago. She is going 40 mph on the way to Chicago and 80 mph on the way back. What is the average traveling time.
Answer by Alan3354(31538)

A person is traveling to Chicago. She is going 40 mph on the way to Chicago and 80 mph on the way back. What is the average traveling time.
-------------------
d = distance one-way
d = rt
t = d/r
Total time = d/40 + d/80 = 3d/80
-------------
Avg r = 2d/t = 2d/(3d/80) = 160/3 mph